# A highly accurate procedure for computing globally optimal Wannier functions in one-dimensional crystalline insulators

### v3

@@ 3.3 Perturbation analysis in the Fourier domain
Remark 3.9.
We observe that the operator $-\mathbf{\Theta}^{\dagger}(k)\mathbf{S}(k)$ in (61) has a spectrum that decays as $1/m$ asymptotically for any $k\in I_{\mathrm{bz}}$, so it is a compact operator on $\ell^{2}(\mathbb{Z})$. As a result, despite the fact that (61) is infinite-dimensional, (61) is very benign both mathematically and numerically.
@@ 3.4 Analysis of the discontinuities at zone boundaries
$a_{m+1}(-\Omega/2)=e^{\mathrm{i}\phi_{\rm zak}}a_{m}(\Omega/2),\quad m\in\mathbb{Z},$ (68)
where $\phi_{\rm zak}$ is a real number independent of $m$.
@@
where $R$ is the right shift operator defined in (2). Since the eigenvalue $E$ is not degenerate, $\mathbf{y}(-\Omega/2)$ and $R(\mathbf{y}(\Omega/2))$ can only differ by a constant of unity modulus $e^{\mathrm{i}\phi_{\rm zak}}$ for some real $\phi_{\rm zak}$. The component form of $\mathbf{y}(-\Omega/2)=e^{\mathrm{i}\phi_{\rm zak}}R(\mathbf{y}(\Omega/2))$ yields (68). ∎
@@
$=-e^{-\mathrm{i}\phi_{\rm zak}}\mathbf{\Theta}^{\dagger}(-\Omega/2)\mathbf{S}(-\Omega/2)\mathbf{y}(-\Omega/2)$ (74)
$=e^{-\mathrm{i}\phi_{\rm zak}}\frac{\mathrm{d}}{\mathrm{d}k}\mathbf{y}(-\Omega/2).$ (75)
This shows that the jump between $a^{\prime}_{m}(\Omega/2)$ and $a^{\prime}_{m+1}(-\Omega/2)$ is also a factor of $e^{i\phi_{\rm zak}}$. Repeated differentiation of (72) and (73) shows the jump between the $n$th derivatives $a^{(n)}_{m}(\Omega/2)$ and $a^{(n)}_{m+1}(-\Omega/2)$ are all identical.
Remark 3.11.
The quantity $\phi_{\rm zak}$ was introduced by Zak in [22] in a different context and is known as the Zak phase. It was shown in [22] that the Zak phase depends only on the Schrödinger equation (9) and determines the center of the Wannier function.
@@ 3.5 Gauge transformation
$\widetilde{\mathbf{y}}(k)=e^{-\mathrm{i}\phi(k)}\mathbf{y}(k)\,,\quad\mbox{ for $k\in I_{\mathrm{bz}}$ },$ (76)
@@
$\widetilde{\alpha}(k+m\Omega)=e^{-\mathrm{i}\phi(k)}\mathbf{y}_{m}(k)=e^{-\mathrm{i}\phi(k)}a_{m}(k),\quad m\in\mathbb{Z},\,k\in I_{\mathrm{bz}}.$ (77)
@@
$\widetilde{\mathbf{y}}^{\prime}(k)=-\mathbf{\Theta}^{\dagger}(k)\mathbf{S}(k)\widetilde{\mathbf{y}}(k)-\mathrm{i}\phi^{\prime}(k)\widetilde{\mathbf{y}}(k),\quad k\in I_{\mathrm{bz}}.$ (78)
In the physics literature, $\phi^{\prime}$ is known as the Berry connection and the phase $\phi$ is known as the Berry phase. The Berry connection is usually denoted by $A$ and defined as $A(k)=\mathrm{i}\innerproduct{u_{k}}{u^{\prime}_{k}}$ in Bra-Ket notation for some function $u_{k}$ satisfying (13). In this paper, this corresponds to the identity:
$\mathrm{i}\widetilde{\mathbf{y}}^{*}(k)\widetilde{\mathbf{y}}^{\prime}(k)=\phi^{\prime}(k)/a,$ (79)
@@
We choose $\phi$ such that $\phi(k)=-\phi(-k)$ and $\phi^{\prime}(k)=\phi_{\rm zak}/\Omega$, where $\phi_{\rm zak}$ is defined in (68):
$\phi(k)=\int_{-\Omega/2}^{k}\phi^{\prime}(k^{\prime})\mathrm{d}k^{\prime}-\frac{\Omega}{2}=\frac{\phi_{\rm zak}}{\Omega}k,\quad k\in I_{\mathrm{bz}}.$ (80)
By Lemma 3.10, this choice of $\phi(k)$ fixes the jump so that
@@
$\widetilde{\mathbf{y}}^{\prime}(-\Omega/2)=-\mathbf{\Theta}^{\dagger}(-\Omega/2)\mathbf{S}(-\Omega/2)\widetilde{\mathbf{y}}(-\Omega/2)-\mathrm{i}\phi^{\prime}(-\Omega/2)\widetilde{\mathbf{y}}(-\Omega/2),$ (82)
$R(\widetilde{\mathbf{y}}^{\prime}(\Omega/2))=-\mathbf{\Theta}^{\dagger}(-\Omega/2)\mathbf{S}(-\Omega/2)R(\widetilde{\mathbf{y}}(\Omega/2))-\mathrm{i}\phi^{\prime}(\Omega/2)R(\widetilde{\mathbf{y}}(\Omega/2)).$ (83)
Since $\phi^{\prime}$ in (80) is a constant, we have $\phi^{\prime}(-\Omega/2)=\phi^{\prime}(\Omega/2)$. Together with (81), we conclude that
@@
We also observe that the choice $\phi^{\prime}(k)=\phi_{\rm zak}/\Omega$ is not the only possibility that yields exponential localization. The above will still hold if $\phi^{\prime}(k)$ is chosen to be a real, periodically analytic function in $I_{\mathrm{bz}}$ with the zeroth Fourier coefficient $(\phi_{\rm zak}+2\pi n)/\Omega$ for any $n\in\mathbb{Z}$. Such functions can be written as
$\phi^{\prime}(k)=\frac{\phi_{\rm zak}+2\pi n}{\Omega}+\sum_{\begin{subarray}{c}m=-\infty\\ m\neq 0\end{subarray}}^{\infty}c_{m}e^{\mathrm{i}mak},$ (86)
where the Fourier coefficients $c_{m}$ decay exponentially as |$\absolutevalue{m}\to\infty$ | → ∞, and the Berry phase $\phi$ is then given by
$\phi(k)=\int_{-\Omega/2}^{k}\phi^{\prime}(k^{\prime})\mathrm{d}k^{\prime}-\frac{\Omega}{2},\quad k\in I_{\mathrm{bz}}.$ (87)
We summarize this fact in the following theorem. It will be the key tool in constructing exponentially localized Wannier functions.
@@
Suppose that $\mathbf{y}(k)$ with $k\in I_{\mathrm{bz}}$ is a solution to (61) with initial conditions (63). Then, $\mathbf{y}(k)$ defines a function $\alpha$ in $\mathbb{R}$ by (54). Suppose further that we choose $\phi^{\prime}(k)$ to be of the form in (86) so that a new function $\widetilde{\mathbf{y}}$ is defined by the gauge transform (76). The function $\widetilde{\alpha}$ defined by $\widetilde{\mathbf{y}}$ in (77) is analytic in $\mathbb{R}$. As a result, the Wannier function as the Fourier transform of $\widetilde{\alpha}$ in (48) is exponentially localized.
@@ 3.6 Gauge choice and Wannier localization
In the previous section, we have seen that many choices of the Berry phase $\phi$ (see Theorem 3.12 and (86)) are able to produce exponentially localized Wannier functions. In this section, we derive the optimal choice in terms of minimizing the variance defined by (52). It is a standard measure of the localization of Wannier functions [15]. First, we modify the formulas in Lemma 52 to account for the inclusion of $\phi$. The following is a result of (77), (78), (79) with $\phi^{\prime}$ given in (86) and the formulas in Lemma 52.
@@
$\langle x\rangle=\frac{a}{2\pi}\left[\phi\left(\frac{\Omega}{2}\right)-\phi\left(-\frac{\Omega}{2}\right)\right]=\frac{\phi_{\rm zak}}{2\pi}a+na,$ (88)
$\langle x^{2}\rangle$ $=$ $\frac{a^{2}}{2\pi}\int_{-\Omega/2}^{\Omega/2}\norm{\mathbf{y}^{\prime}(k)}^{2}\,{\rm d}k+\frac{a}{2\pi}\int_{-\Omega/2}^{\Omega/2}\phi^{\prime}(k)^{2}\,{\rm d}k$ (89)
$=$ $\frac{a^{2}}{2\pi}\int_{-\Omega/2}^{\Omega/2}\norm{\mathbf{y}^{\prime}(k)}^{2}\,{\rm d}k+\left(\frac{\phi_{\rm zak}}{2\pi}a+na\right)^{2}+\sum_{\begin{subarray}{c}m=-\infty\\ m\neq 0\end{subarray}}^{\infty}\absolutevalue{c_{m}}^{2},$ (90)
@@
$\langle x^{2}\rangle-\langle x\rangle^{2}=\frac{a^{2}}{2\pi}\int_{-\Omega/2}^{\Omega/2}\norm{\mathbf{y}^{\prime}(k)}^{2}\,{\rm d}k+\sum_{\begin{subarray}{c}m=-\infty\\ m\neq 0\end{subarray}}^{\infty}\absolutevalue{c_{m}}^{2}.$ (91)
@@
$\phi^{\prime}(k)=\frac{\phi_{\rm zak}+2\pi n}{\Omega},$ (92)
@@
$\phi(k)=\int_{-\Omega/2}^{k}\phi^{\prime}(k^{\prime})\,\mathrm{d}k^{\prime}-\frac{\Omega}{2}=\frac{\phi_{\rm zak}}{\Omega}k,\quad k\in I_{\mathrm{bz}}.$ (93)
@@
$\langle x\rangle=\frac{\phi_{\rm zak}}{2\pi}a.$ (94)
@@
Remark 3.15.
The condition in (92) can also be obtained by applying the calculus of variations on the variance in (91). Since the center in (91) must equal $\frac{\phi_{\rm zak}}{2\pi}a+na$ for (81) to hold, it can be excluded from the objective function. Therefore, to minimize the variance it is sufficient for $\phi$ to minimize $\langle x^{2}\rangle$, where the only gauge-dependent term is
$\int_{-\Omega/2}^{\Omega/2}\phi^{\prime}(k)^{2}\,{\rm d}k.$ (96)
@@
$\phi^{\prime\prime}=0,\quad\mbox{subject to $\phi(\Omega/2)-\phi(-\Omega/2)=\frac{\phi_{\rm zak}}{2\pi}a+na$.}$ (97)
The derivative of the solution is precisely (92). Results of this form are known in physics literature [15, 20, 21], but they are usually stated in a nonconstructive form and are only meant to be theoretical tools characterizing the minimum.
@@ 3.7 Realty of Wannier functions
While we have completed the construction of optimally localized Wannier functions, there is one remaining degree of freedom, the phase choice of the vector $\mathbf{y}_{0}$ in (63); we could replace $\mathbf{y}_{0}$ by $e^{-\mathrm{i}\phi_{0}}\mathbf{y}_{0}$ for any real $\phi_{0}$ and all above formulas would still hold, including the expressions in Corollary 3.14. However, the phase factor $e^{-\mathrm{i}\phi_{0}}$ does change the resulting $W_{0}$. In this section, we prove that the constructed Wannier function $W_{0}$ can always be chosen to be real if $\mathbf{y}_{0}$ in (63) satisfies a simple condition.
@@
Suppose the vector $\widetilde{\mathbf{y}}(k)$ for $k\in I_{\mathrm{bz}}$ and its corresponding function $\widetilde{\alpha}$ are constructed based on Theorem 3.12 with $\phi$ given in Corollary 3.14. Suppose further that the vector $\mathbf{y}(k)=\left\{a_{m}(k)\right\}_{m=-\infty}^{\infty}$ is related to $\widetilde{\mathbf{y}}(k)$ by (76). It is always possible to choose the elements in the initial vector $\mathbf{y}(-\Omega/2)=\mathbf{y}_{0}$ in (63) to satisfy
@@
Replacing the original $\mathbf{y}_{0}$ with $e^{-\mathrm{i}\phi_{0}}\mathbf{y}_{0}$ with $\phi_{0}=\theta/2$ produces $\mathbf{y}$ that satisfies (102).
@@
Together with (77) and $\phi(k)=-\phi(-k)$ in (93), the above relation yields the desired result:
$\widetilde{\alpha}(k+m\Omega)=e^{-\mathrm{i}\phi(k)}a_{m}(k)=e^{\mathrm{i}\phi(-k)}\overline{a}_{-m}(-k)=\overline{\widetilde{\alpha}}(-k-m\Omega)\quad m\in\mathbb{Z}.$ (107)
We observe that the condition in (100) is applicable after solving (61) for determining $\mathbf{y}$ in $[-\Omega/2,\Omega/2]$; we can obtain a constant phase factor $e^{-\mathrm{i}\phi_{0}}$ to enforce (100) so the the resulting $W_{0}$ is real.

### v4

@@ 3.3 Perturbation analysis in the Fourier domain
We observe that the operator $\mathbf{\Theta}^{\dagger}(k)\mathbf{S}(k)$ in (61) has a spectrum that decays as $1/m$ asymptotically for any $k\in I_{\mathrm{bz}}$, so it is a compact operator on $\ell^{2}(\mathbb{Z})$. As a result, despite the fact that (61) is infinite-dimensional, (61) is very benign both mathematically and numerically.
@@ 3.4 Analysis of the discontinuities at zone boundaries
$a_{m}(\Omega/2)=e^{\mathrm{i}\varphi_{\rm zak}}a_{m+1}(-\Omega/2),\quad m\in\mathbb{Z},$ (68)
where $\varphi_{\rm zak}$ is a real number independent of $m$.
@@
where $R$ is the right shift operator defined in (2). Since the eigenvalue $E$ is not degenerate, $\mathbf{y}(-\Omega/2)$ and $R(\mathbf{y}(\Omega/2))$ can only differ by a constant of unity modulus $e^{\mathrm{i}\varphi_{\rm zak}}$ for some real $\varphi_{\rm zak}$. The component form of $e^{\mathrm{i}\varphi_{\rm zak}}\mathbf{y}(-\Omega/2)=R(\mathbf{y}(\Omega/2))$ yields (68). ∎
@@
$=-e^{\mathrm{i}\varphi_{\rm zak}}\mathbf{\Theta}^{\dagger}(-\Omega/2)\mathbf{S}(-\Omega/2)\mathbf{y}(-\Omega/2)$ (74)
$=e^{\mathrm{i}\varphi_{\rm zak}}\frac{\mathrm{d}}{\mathrm{d}k}\mathbf{y}(-\Omega/2).$ (75)
This shows that the jump between $a^{\prime}_{m}(\Omega/2)$ and $a^{\prime}_{m+1}(-\Omega/2)$ is also a factor of $e^{\mathrm{i}\varphi_{\rm zak}}$. Repeated differentiation of (72) and (73) shows the jump between the $n$th derivatives $a^{(n)}_{m}(\Omega/2)$ and $a^{(n)}_{m+1}(-\Omega/2)$ are all identical. The quantity $\varphi_{\rm zak}$ was introduced by Zak in [23] in a different context and is known as the Zak phase. It was shown in [23] that the Zak phase depends only on the Schrödinger equation (9) and determines the center of the Wannier function.
@@ 3.5 Gauge transformation
[…]
@@
$\widetilde{\alpha}(k+m\Omega)=e^{-\mathrm{i}\varphi(k)}\mathbf{y}_{m}(k)=e^{-\mathrm{i}\varphi(k)}a_{m}(k),\quad m\in\mathbb{Z},\,k\in I_{\mathrm{bz}}.$ (77)
@@
$\widetilde{\mathbf{y}}^{\prime}(k)=-\mathbf{\Theta}^{\dagger}(k)\mathbf{S}(k)\widetilde{\mathbf{y}}(k)-\mathrm{i}\varphi^{\prime}(k)\widetilde{\mathbf{y}}(k),\quad k\in I_{\mathrm{bz}}.$ (78)
[…]
$\mathrm{i}\widetilde{\mathbf{y}}^{*}(k)\widetilde{\mathbf{y}}^{\prime}(k)=\varphi^{\prime}(k)/a,$ (79)
@@
We choose $\varphi$ such that $\varphi(k)=-\varphi(-k)$ and $\varphi^{\prime}(k)=\varphi_{\rm zak}/\Omega$, where $\varphi_{\rm zak}$ is defined in (68):
$\varphi(k)=\int_{-\Omega/2}^{k}\varphi^{\prime}(k^{\prime})\mathrm{d}k^{\prime}-\frac{\Omega}{2}=\frac{\varphi_{\rm zak}}{\Omega}k,\quad k\in I_{\mathrm{bz}}.$ (80)
By Lemma 3.9, this choice of $\varphi(k)$ fixes the jump so that
@@
$\widetilde{\mathbf{y}}^{\prime}(-\Omega/2)=-\mathbf{\Theta}^{\dagger}(-\Omega/2)\mathbf{S}(-\Omega/2)\widetilde{\mathbf{y}}(-\Omega/2)-\mathrm{i}\varphi^{\prime}(-\Omega/2)\widetilde{\mathbf{y}}(-\Omega/2),$ (82)
$R(\widetilde{\mathbf{y}}^{\prime}(\Omega/2))=-\mathbf{\Theta}^{\dagger}(-\Omega/2)\mathbf{S}(-\Omega/2)R(\widetilde{\mathbf{y}}(\Omega/2))-\mathrm{i}\varphi^{\prime}(\Omega/2)R(\widetilde{\mathbf{y}}(\Omega/2)).$ (83)
Since $\varphi^{\prime}$ in (80) is a constant, we have $\varphi^{\prime}(-\Omega/2)=\varphi^{\prime}(\Omega/2)$. Together with (81), we conclude that
@@
We also observe that the choice $\varphi^{\prime}(k)=\varphi_{\rm zak}/\Omega$ is not the only possibility that yields exponential localization. The above will still hold if $\varphi^{\prime}$ is chosen to be a real, periodically analytic function on $I_{\mathrm{bz}}$ with the zeroth Fourier coefficient $(\varphi_{\rm zak}+2\pi n)/\Omega$ for any $n\in\mathbb{Z}$. Such functions can be written as
$\varphi^{\prime}(k)=\frac{\varphi_{\rm zak}+2\pi n}{\Omega}+\sum_{\begin{subarray}{c}m=-\infty\\ m\neq 0\end{subarray}}^{\infty}c_{m}e^{\mathrm{i}mak},$ (86)
where the Fourier coefficients $c_{m}$ decay exponentially as $|\absolutevalue{m}\to\infty$, and the Berry phase $\varphi$ is then given by
$\varphi(k)=\int_{-\Omega/2}^{k}\varphi^{\prime}(s)\,\mathrm{d}s-\frac{\Omega}{2},\quad k\in I_{\mathrm{bz}}.$ (87)
We summarize this fact in the following theorem. It will be the key tool for constructing exponentially localized Wannier functions.
@@
Suppose that $\mathbf{y}$ on $I_{\mathrm{bz}}$ is a solution to (61) with initial conditions (63) and $\mathbf{y}$ defines a function $\alpha$ on $\mathbb{R}$ by (54). Suppose further that $\varphi^{\prime}$ is chosen as (86), by which we apply the gauge transform (76) to define a new function $\widetilde{\mathbf{y}}$. Then the function $\widetilde{\alpha}$ defined by $\widetilde{\mathbf{y}}$ in (77) is analytic on $\mathbb{R}$. As a result, the Wannier function as the Fourier transform of $\widetilde{\alpha}$ in (48) is exponentially localized.
@@ 3.6 Gauge choice and Wannier localization
In the previous section, we have seen that many choices of the Berry phase $\varphi$ (see Theorem 3.10 and (86)) are able to produce exponentially localized Wannier functions. In this section, we derive the optimal choice in terms of minimizing the variance defined by (52). It is a standard measure of the localization of Wannier functions [16]. First, we modify the formulas in Lemma 52 to account for the inclusion of $\varphi$. The following is a result of (77), (78), (79) with $\varphi^{\prime}$ given in (86) and the formulas in Lemma 52.
@@
$\langle x\rangle=\frac{a}{2\pi}\left[\varphi\left(\frac{\Omega}{2}\right)-\varphi\left(-\frac{\Omega}{2}\right)\right]=\frac{\varphi_{\rm zak}}{2\pi}a+na,$ (88)
$\langle x^{2}\rangle$ $=$ $\frac{a^{2}}{2\pi}\int_{-\Omega/2}^{\Omega/2}\norm{\mathbf{y}^{\prime}(k)}^{2}\,{\rm d}k+\frac{a}{2\pi}\int_{-\Omega/2}^{\Omega/2}\varphi^{\prime}(k)^{2}\,{\rm d}k$ (89)
[…]
@@
[…]
@@
$\varphi^{\prime}(k)=\frac{\varphi_{\rm zak}+2\pi n}{\Omega},$ (92)
@@
$\varphi(k)=\int_{-\Omega/2}^{k}\varphi^{\prime}(s)\,\mathrm{d}s-\frac{\Omega}{2}=\frac{\varphi_{\rm zak}}{\Omega}k,\quad k\in I_{\mathrm{bz}}.$ (93)
@@
$\langle x\rangle=\frac{\varphi_{\rm zak}}{2\pi}a.$ (94)
@@
We note that the condition in (92) can also be obtained by applying the calculus of variations on the variance in (91). Since the center in (91) must equal $\frac{\varphi_{\rm zak}}{2\pi}a+na$ for (81) to hold, it can be excluded from the objective function. Therefore, to minimize the variance it is sufficient for $\varphi$ to minimize $\langle x^{2}\rangle$, where the only gauge-dependent term is
$\int_{-\Omega/2}^{\Omega/2}\varphi^{\prime}(k)^{2}\,{\rm d}k.$ (96)
@@
$\varphi^{\prime\prime}=0,\quad\mbox{subject to $\varphi(\Omega/2)-\varphi(-\Omega/2)=\frac{\varphi_{\rm zak}}{2\pi}a+na$.}$ (97)
The derivative of the solution is precisely (92). Results of this form are known in physics literature [2, 16].
@@ 3.7 Realty of Wannier functions
While we have completed the construction of optimally localized Wannier functions, there is one remaining degree of freedom, the phase choice of the vector $\mathbf{y}_{0}$ in (63); we could replace $\mathbf{y}_{0}$ by $e^{-\mathrm{i}\varphi_{0}}\mathbf{y}_{0}$ for any real $\varphi_{0}$ and all above formulas would still hold, including the expressions in Corollary 3.12. However, the phase factor $e^{-\mathrm{i}\varphi_{0}}$ does change the resulting $W_{0}$. In this section, we prove that the constructed Wannier function $W_{0}$ can always be chosen to be real if $\mathbf{y}_{0}$ in (63) satisfies a simple condition.
@@
Suppose the vector $\widetilde{\mathbf{y}}(k)$ for $k\in I_{\mathrm{bz}}$ and its corresponding function $\widetilde{\alpha}$ are constructed based on Theorem 3.10 with $\varphi$ given in Corollary 3.12. Suppose further that the vector $\mathbf{y}(k)=\left\{a_{m}(k)\right\}_{m=-\infty}^{\infty}$ is related to $\widetilde{\mathbf{y}}(k)$ by (76). Then it is always possible to choose the elements in the initial vector $\mathbf{y}(-\Omega/2)=\mathbf{y}_{0}$ in (63) to satisfy
@@
Replacing the original $\mathbf{y}_{0}$ with $e^{-\mathrm{i}\varphi_{0}}\mathbf{y}_{0}$ with $\varphi_{0}=\theta/2$ produces $\mathbf{y}$ that satisfies (102).
@@
Together with (77) and $\varphi(k)=-\varphi(-k)$ in (93), the above relation yields the desired result:
$\widetilde{\alpha}(k+m\Omega)=e^{-\mathrm{i}\varphi(k)}a_{m}(k)=e^{\mathrm{i}\varphi(-k)}\overline{a}_{-m}(-k)=\overline{\widetilde{\alpha}}(-k-m\Omega)\quad m\in\mathbb{Z}.$ (107)
We observe that the condition in (100) is applicable after solving (61) for determining $\mathbf{y}$ on $\Omega/2,\Omega/2]$; we can obtain a constant phase factor $e^{-\mathrm{i}\varphi_{0}}$ to enforce (100) so the the resulting $W_{0}$ is real.